For the combustion of cyclopropane at 25 ºC C3H6(g) + (9/2) O2(g) → 3 CO2(g) + 3 H2O(liq) ΔHºc(C3H6,g) = -2091 kJ / mol. What is the standard molar enthalpy of formation of cyclopropane(g) at 25 ºC if the standard molar enthalpies of formation of CO2(g) and H2O(liq) at that temperature are, respectively, -393.51 kJ / mol and -285.83 kJ / mol?
Answer #1

given combustion reaction of cyclopropane is given bellow

C3H6(g) + (9/2)O2  \rightarrow 3CO2 +3H2O \DeltaH0 = -2091KJ/mol

media%2F8f8%2F8f8e01fa-a701-413b-8aa2-d6 H0 C3H6= ? KJ/mol

\DeltarxnH0 = -2091KJ/mol

media%2Fd79%2Fd794d9b3-76b6-49f6-b341-f0fH0 O2(g) = 0 KJ/mol

media%2Fcd3%2Fcd3f780a-ddec-48f6-aa9c-c3fH0 CO2(g) = -393.51 KJ/mol

media%2F8c9%2F8c9e233b-3a30-4644-b060-71fH0 H2O(g) = -285.83 KJ/mol

media%2Fe7a%2Fe7a3286d-993f-40d0-a74d-81 H0 = media%2Fbf7%2Fbf7de54b-d0a5-4eab-b582-2cmedia%2F56c%2F56cd2c1f-aaa0-4ded-bb0a-b2fH0 (Product) - media%2Ff30%2Ff308e28a-f9a2-455f-9db4-a1media%2F346%2F3465fda5-57eb-4b43-9aed-7cfH0 (Reactant)

media%2F927%2F927b4bfe-f2bd-4eef-b66f-65 H0rxn= [(3 media%2Fd33%2Fd33a3000-84d5-4161-937a-e7fH0 CO2) + (3media%2F4fc%2F4fcc5274-fdb5-4496-9263-famedia%2F696%2F696781c7-c2e3-43f8-9038-29fH0 H2O)] - [(media%2F21e%2F21e54d57-577d-45d5-a175-24fH0 C3H6) + (9/2media%2Fc8e%2Fc8e04cb2-c6a7-4214-a13c-4dmedia%2F454%2F454681be-9607-43b4-8dfd-19fH0 O2)]

-2091 = [(3media%2F62f%2F62fd93c9-c428-4d1d-a499-20 (-393.15)) + (3media%2F68b%2F68be0236-1d5f-421a-b251-f6 (-285.83))] - [\DeltafH0C3H9 + (9/2 (0)]

-2091 = [-2036.94] - [\DeltafHC3H9]

media%2Ffcb%2Ffcb2fc5a-e8b6-4b23-a22c-d5fH0C3H9 = 2091 - 2036.94 = +54.06KJ/mol

ans standard molar enthalpy of formation of cyclopropane is + 54.06 KJ/mol